# Joint Distributed Generation Maximization and Radial Distribution Network Reconfiguration

### v1

@@ II-D DG maximizing ROPF Model and Solution Approach
In (12), the decision variables without subscripts denote the vectors or matrices stacked by the corresponding variables for individual buses or lines (e.g., $p^{G}$ is the $N$-vector of all $p^{G}_{n}$ for $n\in\mathcal{N}$). $\nu,\ell,p^{G},q^{G},P,Q$ are continuous vector variables, while $\pi$ and $\alpha$ are 0-1 binary matrix variables. Except for (5d), all constraints can be rewritten as linear equalities or inequalities (cf. (6)). In contrast, even after linearization (5d) involves bilinear equality constraints because of the product of continuous variables $\nu_{k}$ and $\ell_{n}$. Thus, problem (12) is a mixed integer linear program with additional (typically non-convex) bilinear equality constraints. This problem is more general (but more difficult) than mixed integer linear program which is common in the literature.
Gradient-based local optimization methods (e.g., interior point method) and meta-heuristics (e.g., ant colony optimization) are common solution methods for (12). However, we propose to use the SBB (i.e., spatial branch-and-bound) algorithm. We call the approach to solve (12) using SBB our “proposed method”. Our choice is motivated by the following relative advantages: (a) SBB returns (sub)optimal solutions together with certificates of degree of optimality, (b) SBB directly uses the model and its structure in the search of optimal solutions, and (c) for realistically sized benchmark instances of (12) in Section III and Section IV, the proposed method returns dispatch control decisions with certifiable optimality in a few minutes, making it suitable even for control center operations.
An alternative to (12) is to replace the bilinear equalities in (5d) with inequalities, leading to a relaxation [15] as
subject to (2) to (5) except for (5d), (7) to (9) and
Unlike (5d), the rotated second order cone inequalities in (13b) are easier to handle computationally (e.g., [25, p. 197]). This allows (13) to be categorized as a mixed integer conic program, which becomes a convex problem when all integer variables are fixed (e.g., during branch-and-bound calculations). This is the main advantage of (13). However, a major question regarding (13) is whether it is tight in the sense that (13b) hold as equalities at optimality. If (13) is not tight, the control decisions usually lead to violations of grid constraints (e.g., voltage and/or current limits) as the DistFlow equations are not satisfied. When replacing the objective in (13) with “standard” ones (e.g., total line loss minimization, quadratic fuel cost minimization with nonnegative coefficients, conservation voltage reduction), (13) appears to be tight as demonstrated in [16, 2, 17]. When $K=0$ (i.e., OPF) (13) can be shown to be tight for the standard objectives under some additional assumptions (e.g., [15]). However, with DG maximization as studied in this paper no theoretical guarantee of the tightness of (13) has been shown even for $K=0$. While [19] demonstrated through a numerical case study that a version of (13) with more operational considerations delivered sensible control decisions (tightness of (13) was not examined in [19]), in Section III-D we provide a simple three-bus counterexample to show that (13) is not tight and the resulting control will lead to grid constraint violations (this also agrees with the observations in [26]). Hence, tightness of (13) should not be taken for granted and similar extensions of conic relaxation should be studied with care. This motivates us to keep (5d) as equalities in (12) in our proposed method. The computation aspects of (12) will be discussed in greater detail in Section III.
III Illustration and Justification
To justify necessity, we compare the proposed method with potential alternatives to solve the ROPF problem (12) for distribution system benchmarks listed in Table I.
TABLE I: Distribution system benchmark data
@@ III Illustration and Justification
The topology, line parameters and load data of the benchmarks can be found in the respective sources in Table I. For each benchmark we add DG sources in the system (location details omitted). The rating of each DG source is 10 MVA for all benchmarks except for REDS 873+7 where the rating is 50 MVA (large enough to prevent the capacity constraint (7b) from being active). We note that 136ma and 135+1 are in fact the same network with different DG locations.
@@
All computation in this section is performed on a PC with a 24-core Intel i9-12900K CPU at 3.2 GHz and 64 GB of RAM. The ROPF problem in (12) is solved with Gurobi 10.0.1 (with parameter “non-convex” set to 2) and we use AMPL to implement the optimization models. Gurobi is called with its default parameter values (e.g., target optimality gap 0.01%).
III-A Reconfiguration with the 33-bus Benchmark
We consider the 33-bus benchmark 33bw described in Table I. The topology, the locations of the two DG sources and the line current ratings are shown in Fig. 1.
@@ III-A Reconfiguration with the 33-bus Benchmark
We first attempt to solve the reconfiguration problem in (12) maximizing the total DG output with $K=0$ (i.e., OPF with original network topology). The instance is infeasible due to voltage restrictions. However, if we relax the voltage lower limit from 0.95 pu to 0.9 pu, the instance becomes feasible and the voltages from bus 12 to bus 17 (towards the end of the feeder) are below 0.95 pu. Next, we solve (12) with $K=2$ (i.e., one switch opened and another switch closed) with the usual voltage limits of 0.95 pu and 1.05 pu. The result is shown in Fig. 2, with a total DG output of 4.33 MW.
@@
The aforementioned results help us to identify the segment between bus 5 and bus 28 to be the bottleneck. If we increase the current ratings of these four lines (e.g., due to dynamic line rating), we expect the DG output to further increase. Indeed, if the ratings are increased from 200 A to 300 A, then for $K=0,2,4,6$ the total DG outputs are, respectively, 0 MW (infeasible instance), 5.44 MW, 6.68 MW, and 6.77 MW.
III-B Computation Time Evaluation
The elapsed runtime to solve the benchmark instances is summarized in Table II.
TABLE II: Runtime [s] to achieve 0.01% optimality gap, or optimality gap [%] after one-hour runtime, for benchmark instances of (12)
@@ III-B Computation Time Evaluation
In the table, N/A means that the instance is infeasible. In addition, for REDS 135+1 with $K=8$ and REDS 873+7 with $K=6$ and $K=8$ the table entries show the percentage optimality gap after time limit of one hour. In general, runtime increases as the network size and $K$ increase. However, there are exceptions as REDS 135+1 appears to be difficult for $K=8$ (cf. 136ma with $K=8$ differing only in DG locations), whereas REDS 83+11 is easy (even with $K=26$ the runtime is 11.8s, not shown in Table II). Despite some irregularities in the runtime pattern, we conclude that the proposed method can reliably obtain guaranteed optimal DG set points and network configuration for moderately-sized instances (e.g., networks up to few hundred buses and $K\leq 6$). However, discretion is needed when attempting to solve larger instances. Nevertheless, typically the marginal benefit diminishes rapidly with increasing $K$ [2]. Furthermore, despite not being able to solve large instances to optimality, the proposed method can return suboptimal solutions with certified optimality gap. Runtime to achieve 5% optimality gap for the benchmarks is summarized in Table III.
TABLE III: Runtime to achieve 5% optimality gap for benchmark instances of (12)
@@
III-C Comparison with Local Optimization Solvers
In this part of case study, we compare the OPF results by solving (12) with $K=0$ and by using Matpower 7.1’s OPF subroutine runopf.m. The Matpower OPF formulation is based on the AC power flow equations (voltage phasors in polar coordinates) instead of the DistFlow equations in (1) in the proposed model. To maximize the total DG output, the Matpower OPF cost function is set to $\sum_{n}c^{0}_{n}+c^{1}_{n}p_{n}^{G}+c^{2}_{n}(p_{n}^{G})^{2}$ with $c^{0}_{n}=c^{2}_{n}=0$ and $c^{1}_{n}=-1$ for all $n\in\mathcal{N}$. This is not the typical positive quadratic cost function. We run Matpower with the default options (e.g., flat start initial guess) and with two optimization solver choices: (a) MIPS interior point method and (b) Matlab’s fmincon active set method. For the feasible instances in Table II, Matpower OPF solves the benchmark OPF instances with relative ease except that MIPS fails to converge for the 201+3 case. However, Matpower OPF does not provide the optimality guarantee available in the proposed method. Furthermore, Matpower OPF results are sensitive to problem instance data and algorithmic choices. For example, if we increase line current ratings to 600 A for all lines and slightly decrease the DG ratings to 8 MW, the OPF instance for the 33bw benchmark becomes feasible but the Matpower OPF results vary greatly depending on the solver choice and initial guess strategy. Table IV shows the DG maximization results, where the Matpower initial guess strategies include: 0 for flat start, 1 for ignoring the system state (i.e., voltages and generation) in the Matpower case, 2a and 2b for using the system state in the Matpower case (without and with update by first solving (12)) and 3a and 3b for solving power flow and using the resulting state (without and with update by first solving (12)). Table IV indicates that local optimization methods implemented in Matpower cannot reliably solve the DG maximization problem. On the other hand, the proposed method yields the desired DG value of 10.45 MW. Further, we note that the difficulty exhibited in Table IV is intrinsic to the DG maximization objective (as in (12)). If we instead minimize $\sum_{n}c^{0}_{n}+c^{1}_{n}p_{n}^{G}+c^{2}_{n}(p_{n}^{G})^{2}$ with $c^{0}_{n}=c^{2}_{n}=0$ and $c^{1}_{n}=1$ (instead of $-1$) for all $n$, Matpower can reliably obtain the minimum total DG of 1.93 MW using both MIPS and fmincon for all choices of initial guess strategies.
TABLE IV: 33bw 600 A, Matpower OPF, max DG output
@@ III-C Comparison with Local Optimization Solvers
III-D Comparing Mixed Integer Conic Programming Relaxation
In this part of case study, we evaluate the mixed integer conic relaxation approach in (13). A central question regarding the use of (13) is whether (13) is tight (i.e., (13b) are satisfied as equalities at optimality). Here we present a counterexample: Consider the OPF problem for the three-bus example in Fig. 4.
Figure 4: 3-bus distribution system example demonstrating failure of the relaxation approach. The voltage at bus 0 is 1 pu.
We use Gurobi to solve (12) and (13) with $K=0$ (using the same AMPL code except for the difference between (5d) and (13b) and the Gurobi “non-convex” option setting). The results are summarized in Table V, showing different values of DG outputs highlighted in red (indicating failure of conic relaxation).
TABLE V: Per unit OPF results for Fig. 4 by (12) and (13)

### v2

@@ 2.4 DG maximizing ROPF Model and the Proposed Solution Approach
In (12), the decision variables without subscripts denote the vectors or matrices stacked by the corresponding variables for individual buses or lines (e.g., $p^{G}$ is the NN-vector of all $p^{G}_{n}$ for $n\in\mathcal{N}$). $\nu,\ell,p^{G},q^{G},P,Q$ are continuous vector variables, while $\pi$ and $\alpha$ are 0-1 binary matrix variables. Except for (5d), all constraints can be rewritten as linear equalities or inequalities (cf. (6)). In contrast, in (5d) the product of continuous variables $\nu_{k}$ and $\ell_{n}$ cannot be equivalently replaced by any auxiliary variable as in (6) (the transformation requires at least one 0-1 binary variable in the product). Thus, problem (12) is a mixed integer linear program with additional non-convex bilinear equality constraints.
Even though (12) resembles a classical model of the ROPF problem (if the objective is changed to loss minimization), we emphasize that the DG maximization objective in (12) leads to significant ramifications that render most of the typical solution methods ineffective. The details will be provided in Section 4. Instead, we propose to solve (12) using the SBB (i.e., spatial branch-and-bound) algorithm of Gurobi (version 9 or later), and we subsequently call this our “proposed method”. The rationale of our choice is as follows:
  1. (a)
    SBB returns (sub)optimal solutions with certifiable optimality gap.
  2. (b)
    The SBB implementation in Gurobi is specialized in bilinear programming problem, allowing more effective use of the problem structure.
  3. (c)
    Case studies with realistically sized benchmark instances indicate the practical computational performance of the proposed method.
We note that different packages implementing the SBB algorithm (e.g., Gurobi, BARON, Knitro) behave differently for problem (12). The numerical experiment in Section 4.5 indicates that Gurobi is the best performer among our choices.
3 Illustration and Discussion of the Proposed Method
We evaluate the performance of the proposed method by solving instances of the ROPF problem in (12) derived from benchmarks listed in Table 1.
Table 1: Distribution system benchmark data
@@ 3 Illustration and Discussion of the Proposed Method
The topology, line parameters and load data of the benchmarks can be found in the respective sources in Table 1. For each benchmark we add DG sources in the system (location details omitted). The rating of each DG source is 10 MVA for all benchmarks except for REDS 873+7 where the rating is 50 MVA (large enough to prevent the capacity constraint (7b) from being active).
@@
All computation in this and the next section is performed on a PC with an Intel i9-12900K CPU and 64 GB of RAM. The ROPF problem in (12) is solved with Gurobi 10.0.1 (with parameter “non-convex” set to 2) and we use AMPL to implement the optimization models. Gurobi is called with its default parameter values (e.g., target optimality gap 0.01%).
3.1 Reconfiguration with the 33-bus Benchmark
We consider the 33-bus benchmark 33bw described in Table 1. The topology, the locations of the two DG sources and the line current ratings are shown in Fig. 1.
@@ 3.1 Reconfiguration with the 33-bus Benchmark
We first attempt to solve the reconfiguration problem in (12) maximizing the total DG output with $K=0$ (i.e., OPF with original network topology). The instance is infeasible due to operational constraints. However, if we relax the voltage lower limit from 0.95 pu to 0.9 pu, the instance becomes feasible and the voltages from bus 12 to bus 17 (towards the end of the feeder) are below 0.95 pu. Next, we solve (12) with $K=2$ (i.e., one switch opened and another switch closed) with the usual voltage limits of 0.95 pu and 1.05 pu. The result is shown in Fig. 2, with a total DG output of 4.33 MW.
@@
3.2 Computation Time Evaluation
The elapsed runtime to solve the benchmark instances is summarized in Table 2.
Table 2: Runtime [s] to achieve 0.01% optimality gap, or optimality gap [%] after one-hour runtime, for benchmark instances of (12)
@@ 3.2 Computation Time Evaluation
In the table, N/A means that the instance is infeasible. In addition, for REDS 135+1 with $K=8$ and REDS 873+7 with $K=6$ and $K=8$ the table entries show the percentage optimality gap after time limit of one hour. In general, runtime increases as the network size and $K$ increase. However, there are exceptions as REDS 135+1 appears to be difficult for $K=8$ (cf. 136ma with $K=8$ differing only in DG settings), whereas REDS 83+11 is easy (even with $K=26$ the runtime is 11.8s, not shown in Table 2). Despite some irregularities in the runtime pattern, we conclude that the proposed method can reliably obtain guaranteed optimal DG set points and network configuration for moderately-sized instances (e.g., networks up to few hundred buses and $K\leq 6$). However, discretion is needed when solving larger instances. Nevertheless, typically the marginal benefit diminishes rapidly with increasing $K$ (e.g., [15]). Further, despite not being able to solve large instances to optimality, the proposed method can return suboptimal solutions with certified optimality gap. Runtime to achieve 5% optimality gap for the benchmarks is summarized in Table 3.
Table 3: Runtime to achieve 5% optimality gap for benchmark instances of (12)
@@
We conclude that, with reasonable target of optimality gap, the proposed method for DG maximizing ROPF is practical for networks of reasonable size and number of switch status changes. However, for very large number of switch status changes (or a problem to plan for possible line expansion), the proposed method should be augmented with more efficient methodologies.
4 On the Choice of Optimization Solvers and Problem Formulations
This section justifies the choice of our proposed method. We demonstrate that alternative methods for DG maximizing ROPF may not work as intended. These include interior point algorithm, MICP relaxation, linearized DistFlow model, alternative power flow equation formulations and SBB implementations.
4.1 Pitfalls with Local Optimization Solvers
We evaluate MATPOWER’s OPF subroutine runopf.m to solve a special case of (12) with $K=0$. To maximize the total DG output, MATPOWER is set to minimize $\sum_{n}c^{0}_{n}+c^{1}_{n}p_{n}^{G}+c^{2}_{n}(p_{n}^{G})^{2}$ with $c^{0}_{n}=c^{2}_{n}=0$ and $c^{1}_{n}=-1$ for all $n\in\mathcal{N}$. This is not the typical positive quadratic cost function. We run MATPOWER with the default options (e.g., flat start initial guess) and with two optimization solver choices: (a) MIPS interior point method and (b) Matlab’s fmincon active set method. For the feasible instances in Table 2, MATPOWER solves the benchmark OPF instances with relative ease except that MIPS fails to converge for the 201+3 case. However, MATPOWER results are sensitive to problem instance data and algorithmic choices. For example, for benchmark 33bw with line current ratings increased to 600 A for all lines and DG ratings decreased to 8 MW, the OPF instance becomes feasible but MATPOWER results vary greatly depending on the solver choice and initial guess strategy. Table 4 shows the DG maximization results due to the following initial guess strategies [36]: 0 for flat start, 1 for ignoring the system state in the MATPOWER case, 2a and 2b for using the system state in the MATPOWER case (without and with update by first solving (12)) and 3a and 3b for solving power flow and using the resulting state (without and with update by first solving (12)). Table 4 indicates that local optimization methods as in MATPOWER cannot reliably solve the DG maximization problem. On the other hand, the proposed method returns the true optimal DG value of 10.45 MW.
The difficulty exhibited in Table 4 is intrinsic to the DG maximization objective (as in (12)). If we instead minimize $\sum_{n}c^{0}_{n}+c^{1}_{n}p_{n}^{G}+c^{2}_{n}(p_{n}^{G})^{2}$ with $c^{0}_{n}=c^{2}_{n}=0$ and $c^{1}_{n}=1$ (instead of 1-1) for all $n$, MATPOWER can reliably obtain the minimum total DG of 1.93 MW using both MIPS and fmincon for all choices of initial guess strategies.
Table 4: 33bw 600 A, MATPOWER OPF, max DG output
@@ 4.1 Pitfalls with Local Optimization Solvers
Therefore, when a ROPF problem with “non-conventional” objective (e.g., DG maximization) is considered, caution must be exercised regarding whether the optimization solver should work as intended.
4.2 Pitfalls with MICP Relaxation
An alternative to (12) is to replace the bilinear equations in (5d) with inequalities, leading to a relaxation [8] as
subject to (2) to (5) except for (5d), (7) to (9) and
Unlike (5d), inequalities (13b) allow (13) to be categorized as a MICP problem amenable to efficient solution algorithms. However, a major question regarding (13) is whether it is tight in the sense that (13b) hold as equations at optimality. If (13) is not tight, the control decisions may violate grid constraints (e.g., voltage and current limits) as the DistFlow equations are not satisfied. For (13) with “standard” objective such as total line loss minimization, the tightness of (13) has been empirically and theoretically verified [33, 15, 17, 8]. However, for DG maximization, the issue of tightness has not been explored in depth. In fact, our answer to this question is negative: we will present a three-bus counterexample to show that (13) is not tight and the resulting control will lead to grid constraint violations (this also agrees with the observations in [34]). Consider the OPF problem for the three-bus example in Fig. 4.
Figure 4: 3-bus system showing failure of MICP relaxation. The voltage at bus 0 is 1 pu.
We use Gurobi to solve (12) and (13) with $K=0$. The results are summarized in Table 5, demonstrating that the proposed method is able to obtain the correct optimal DG outputs whereas MICP relaxation fails to do so (the different values of DG outputs are highlighted in red).
Table 5: Per unit OPF results for Fig. 4 by (12) and (13)
@@ 4.2 Pitfalls with MICP Relaxation
bus $\nu$ $\ell$ $p^{G}$ $q^{G}$ $P$ $Q$
1 1.1025 25 7.7518 0.39754 4.9991-4.9991 0.092606
2 1.0964 0.2645 0 0 0.50264 0.19736-0.19736
Conic relaxation in (13)
bus $\nu$ $\ell$ $p^{G}$ $q^{G}$ $P$ $Q$
1 1.1025 25 7.9991 0.64489 4.9991-4.9991 0.092606
2 1.0915 25 0 0 0.75 0.049999
With the problem data in Fig. 4 and the variable values in Table 5, we can verify that the solutions of (12) and (13) satisfy all constraints in the respective problems. For example, the reactive power balance equation in (1b) for bus 1 for (12) is $q^{G}_{1}-q^{L}_{1}=0.39754-0.5=-0.10246=(-1)\times 0.092606+(-0.19736)+0.0075\times 25=-Q_{1}+Q_{2}+x_{01}\ell_{1}$. The analogous equation holds for the solution of (13) as well: $q^{G}_{1}-q^{L}_{1}=0.64489-0.5=0.14489=(-1)\times 0.092606+0.049999+0.0075\times 25=-Q_{1}+Q_{2}+x_{01}\ell_{1}$. However, while equation (5d) at bus 2 holds for the solution of (12): $\nu_{1}\ell_{2}=1.1025\times 0.2645=0.2916=0.50264^{2}+{(-0.19736)}^{2}=P_{2}^{2}+Q_{2}^{2}$, (13b) remains an inequality instead of equality for the solution of (13): $\nu_{1}\ell_{2}=1.1025\times 25=27.5625>0.565=0.75^{2}+0.049999^{2}=P_{2}^{2}+Q_{2}^{2}$. Indeed, a posterior load flow analysis based on $(p^{G},q^{G})$ from (13) reveals that the bus voltages are 1, 1.0539 and 1.0511 per unit, respectively. Similarly, the currents for the lines are 522.53 A and 51.235 A. The voltage and current upper limits are both violated by the solution of MICP relaxation. The aforementioned comparison is repeated for all benchmarks in Table 1. MICP relaxation solutions fail to satisfy equality (5d) in all cases, and result in voltage and current violations.
The objective functions in (12) and (13) have a significant impact on the tightness of MICP relaxation. Following [4], we consider the trade-off between DG maximization and line loss minimization by minimizing in (12) and (13) the following weighted sum objective:
where $0\leq\rho\leq 1$ is a trade-off parameter. With $\rho$ increasing from 0 to 1, the minimization of (14) transitions from DG maximization to loss minimization. By solving (12) and (13) with (14) and running a posterior load flow analysis based on the resulting DG set points and network configuration, the trade-off can be visualized. Fig. 5 shows the Pareto curves for the 33bw benchmark with universal line rating of 600 A and $K=6$.
Figure 5: Pareto curves of 33bw benchmark (600A line rating) due to the proposed model in (12) and MICP relaxation in (13)
For $\rho>0.5$, the two Pareto curves coincide since MICP relaxation solution satisfies (5d). However, when $\rho\leq 0.5$ MICP relaxation fails as it no longer depicts the true trade-off. This case study raises the awareness that the MICP relaxation approach should be used with caution especially when non-conventional objective function is considered.
4.3 Pitfalls with Linearized DistFlow Approximation
We evaluate a variant of (12) by replacing the DistFlow equations in (5) with its linear approximation (i.e., LinDistFlow equations [3]). The LinDistFlow variant ignores the loss-related terms multiplying $\ell_{n}$ in (5a), (5b) and (5c). Also, equation (5d) is dropped. Substituting the DistFlow equations with the LinDistFlow equations changes (12) to a MILP, which is relatively easy to solve. However, the voltage variables in the LinDistFlow variant overestimate the true voltages (e.g., [30, eq. (2a) and (3)]). Furthermore, since the squared current term $\ell$ is missing it is impossible to impose current upper limit constraints. Indeed, for our benchmark instances severe violation of current upper limit is experienced. To counter this, a surrogate current upper limit constraint can be imposed, similar to (5d), as
In (15) the product in the left-hand side can be equivalently replaced by linear terms as illustrated in (6). Below we call LinDistFlow the variant of (12) with LinDistFlow equations and (15). The inclusion of (15) relieves some undervoltage and overcurrent issues, but this is not enough. For instance, for the 33bw benchmark with 600 A line rating and $K=2$, the actual voltages by a posterior load flow analysis and the voltages perceived to be true by the LinDistFlow optimization model (i.e., $(\nu)^{1/2}$) are very different as indicated by Fig. 6.
Also, current upper limits are violated in multiple lines (worst case violation being 115% of rating) despite the surrogate constraint in (15). This is likely to trigger an undervoltage trip. With reference to (5d), overestimation of voltages will lead to underestimation of currents. Indeed, current upper limit are violated in multiple lines (worst case violation being 115% of rating) despite the surrogate constraint in (15). For LinDistFlow with $K=4$ the result is even more extreme. The load flow analysis by Matpower fails to converge. In contrast, solving (12) for the 33bw 600 A benchmark with $K=2$ and $K=4$, respectively, reliably leads to constraint-abiding control with guaranteed optimality.
4.4 Effect of Power Flow Equation Models
The model of power system physcis (e.g., DistFlow equations versus AC power flow equations) significantly affects the solver’s performance to solve (12). Here we investigate two ROPF models containing power flow equations other than the DistFlow equations (5) adopted in our model in (12):
  • ACOPFC: ROPF model based on standard AC power flow equations by Capitanescu et. al. in [5].
  • ACOPFJ: ROPF model based on modified AC power flow equations by Jabr et. al. in [15].
ACOPFC follows the model in [5]. On the other hand, ACOPFJ is based on the following power system physics model in [15]:
$p^{G}_{n}-p^{L}_{n}=\sum\limits_{k:(n,k)\in\mathcal{L}}P_{nk},$ $n\in\mathcal{N}$ (16a)
$q^{G}_{n}-q^{L}_{n}=\sum\limits_{k:(n,k)\in\mathcal{L}}Q_{nk},$ $n\in\mathcal{N}$ (16b)
$P_{nk}=\sqrt{2}g_{nk}u^{(n,k)}-g_{nk}R_{nk}-b_{nk}T_{nk},$ $(n,k)\in\mathcal{L}$ (16c)
$Q_{nk}=-\sqrt{2}b_{nk}u^{(n,k)}+b_{nk}R_{nk}-g_{nk}T_{nk},$ $(n,k)\in\mathcal{L}$ (16d)
$2u^{(n,k)}u^{(k,n)}\geq R_{nk}^{2}+T_{nk}^{2},$ $(n,k)\in\mathcal{L}$ (16e)
$R_{nk}=R_{kn},$ $(n,k)\in\mathcal{L}$ (16f)
$T_{nk}=-T_{kn},$ $(n,k)\in\mathcal{L}$ (16g)
$0\leq u^{(n,k)}\leq\frac{\overline{\nu}_{n}}{\sqrt{2}}\alpha_{nk},$ $(n,k)\in\mathcal{L}$ (16h)
$0\leq u_{n}-u^{(n,k)}\leq\frac{\overline{\nu}_{n}}{\sqrt{2}}(1-\alpha_{nk}),$ $(n,k)\in\mathcal{L}$ (16i)
where $\mathcal{L}$ is the set of directed lines (normally open and normally closed) such that if $\{n,k\}$ is a line then $(n,k),(k,n)\in\mathcal{L}$, and $g_{nk}+jb_{nk}$ is the series admittance of line $(n,k)$. The decision variables include $P_{nk}$, $Q_{nk}$, $u^{(n,k)}$, $R_{nk}$, $T_{nk}$ and $u_{n}$ for all $(n,k)\in\mathcal{L}$. The ROPF formulation with (16) imposed is a relaxation of (12) because of the inequalities in (16e). However, replacing (16e) with equalities
$2u^{(n,k)}u^{(k,n)}=R_{nk}^{2}+T_{nk}^{2},\quad(n,k)\in\mathcal{L}$ (17)
makes (16) equivalent to the exact AC power flow equations (also equivalent to our DistFlow equations in (5)). The ROPF model with (16) (where the (16e) is replaced by (17)) is the ACOPFJ formulation considered in our case study.
Both ACOPFC and ACOPFJ can be written as MINLP with bilinear equality constraints that is solvable using Gurobi. We test solving these models for the 33bw benchmark using Gurobi with time limit of 300 seconds. Table 6 shows the best DG output (if a feasible solution is found), the percentage optimality gap and the runtime to achieve the gaps. It is evident that for the DG maximizing ROPF, our proposed model in (12) is the only practical choice when compared with ACOPFC and ACOPFJ.
Table 6: Comparing (12), ACOPFC and ACOPFJ, 33bw
33bw $K=2$ $K=4$ $K=6$ $K=8$
(12) 4.34 MW 5.67 MW 5.72 MW 5.72 MW
0%, 1.56s 0%, 1.69s 0%, 2.21s 0%, 4.06s
ACOPFC N/A N/A N/A 4.36 MW
fail, 300s fail, 300s fail, 300s 359%, 300s
ACOPFJ N/A 4.46 MW 4.60 MW 4.60 MW
fail 174%, 300s 202%, 300s 191%, 300s
4.5 Effect of SBB Solver Packages
Different implementations of the SBB algorithm behave very differently when solving (12). Here we evaluate three representative commercial packages:
  • Gurobi 10.0.1: specialized for MINLP allowing only non-convex bilinear equations (and convex constraints).
  • BARON 24.5.8: general SBB based global MINLP solver.
  • Knitro 14.0.0: convex MINLP solver using SBB and interior point algorithms.
The three packages are run with default algorithmic parameter settings. Table 7 shows the comparison result for the benchmark 33bw, showing how the three packages behave differently. The fastest runtime of Gurobi can be explained by the fact that it is specialized to consider only bilinear nonlinearity, whereas BARON is slowest as it aims to handle a much wider class of nonlinearities. Knitro is slower than Gurobi, and it may not return the true optimum as indicated by the MW results for $K=2$ (4.14 MW versus 4.34 MW).
Table 7: Comparing Gurobi, BARON and Knitro, 33bw
33bw $K=2$ $K=4$ $K=6$ $K=8$
Gurobi 4.34 MW 5.67 MW 5.72 MW 5.72 MW
1.56s 1.69s 2.21s 4.06s
BARON 4.34 MW 5.67 MW 5.72 MW 5.72 MW
30.1s 43.0s 68.2s 96.4s
Knitro 4.14 MW 5.67 MW 5.71 MW 5.72 MW
8.64s 8.68s 16.5s 49.5s
The result of a similar comparison case study with the 118zh benchmark is shown in Table 8, also indicating Gurobi’s robustness for the purpose of solving (12).
Table 8: Comparing Gurobi, BARON and Knitro, 118zh
118zh $K=0$ $K=2$ $K=4$ $K=6$
Gurobi 27.47 MW 28.57 MW 29.44 MW 30.17 MW
0.82s 20.2s 28.7s 105.5s
BARON 27.47 MW 28.56 MW 29.27 MW 30.15 MW
2.10s 570s 2964s 2hr (time limit)
Knitro 27.46 MW 28.18 MW 29.35 MW 29.49 MW
0.41s 183s 1034s 1176s
5 Practical Case Study - Real Network
5.1 533-bus Distribution System Description
In this section, the ROPF problem (12) is applied to a real medium voltage distribution system in southern Sweden: the 533-bus network described in [21]. With the expansion of DG, many DSOs wish to assess to what extent DG can be integrated in their systems without violating operational limits. This is referred to as performing a hosting capacity (HC) analysis. In [14], four operational limits that constrain the HC in distribution networks are outlined – voltage, overloading, protection and harmonics. With our proposed method we are imposing voltage and current limits, thereby dealing with the first two. Protection is partly included by requiring radial operation, but no further protection limits (i.e. short-circuit current) are set up. The system is assumed to be three-phase balanced without overtones, thus harmonics mitigation is not considered. Within these limits, the HC can be assessed and enhanced, using e.g. network reconfiguration, and our proposed ROPF method does this jointly.